OpenStudy (keeli_grace):

Can y'all help me add and simplify this equation? (: (2x^2 – 11x) + (3x^2 + 10x – 4)

OpenStudy (keeli_grace):

@epoweritheta

OpenStudy (epoweritheta):

add the like terms and say

OpenStudy (keeli_grace):

So, 5x^4 - 4?

OpenStudy (epoweritheta):

see you cant add the exponents , see i will solve this , 5x^2-x-4

OpenStudy (anonymous):

do the terms that are the same ie x^2 and then x: 2x^2 + 3x^2 = 5x^2 -11x + 10x = -x then join them up and put the constant on the end: 5x^2 -x +4

OpenStudy (anonymous):

sorry -4*

OpenStudy (keeli_grace):

Thank you guys! That was amazingly helpful!!! :)

OpenStudy (epoweritheta):

ya -4 sorry @ben1 thank you

OpenStudy (keeli_grace):

Do you think you two could help me with a couple more problems like this? @ben1 and @epoweritheta

OpenStudy (epoweritheta):

ya ....

OpenStudy (keeli_grace):

(3x^2 – 6x + 7) – (x^2 – 6x + 4)

OpenStudy (keeli_grace):

Thank you, this equation needs to be subtracted and simplified, how would I go about doing that?

OpenStudy (epoweritheta):

see @Keeli_Grace do the same thing here look at the previous proble,m and see what i have done there

OpenStudy (keeli_grace):

Okay, one sec. :)

OpenStudy (keeli_grace):

3x^2 + x^2 -6x + -6x 7+4 Is this how I would start?

OpenStudy (keeli_grace):

Wait, no. It's only subtraction, so I messed up somewhere.

OpenStudy (epoweritheta):

ya check again

OpenStudy (keeli_grace):

3x^2 - x^2 = 2x^2 6x - 6x = 0 7-4 = 3 So would it be: 2x^2 + 3?

OpenStudy (epoweritheta):

You are fantastic , you got it

OpenStudy (keeli_grace):

Seriously?!?! Wow! Sweet! Thank you! =D

OpenStudy (epoweritheta):

ya...